In 2018, the Westgate Construction Company entered into a contract to construct a road for Santa Clara County for $10,000,000. The road was completed in 2020. Information related to the contract is as follows:

2018

2019

2020

Cost incurred during the year

$

2,490,000

$

3,984,000

$

2,008,600

Estimated costs to complete as of year-end

5,810,000

1,826,000

0

Billings during the year

2,030,000

4,444,000

3,526,000

Cash collections during the year

1,815,000

3,900,000

4,285,000


Westgate recognizes revenue over time according to percentage of completion.


3. Complete the information required below to prepare a partial balance sheet for 2018 and 2019 showing any items related to the contract. (Do not round intermediate calculations.)

Balance Sheet (Partial)

2018

2019

Current assets:

Accounts receivable

$215,000

Construction in progress

Less: Billings

(2,030,000)

Costs and profit in excess of billings

(2,030,000)

0

Current liabilities:

2019 is cut off, I need that as well.
Answer #1

Calculations of the percentage of completion:

All amounts are in $

Particulars 2018 2019 2020
Cost Incurred during the year (a) 2490000 3984000 2008600
Accumulated Costs (b) 2490000 6474000 8482600
Further Cost Estimated to complete the project (c) 5810000 1826000 0
Total (d=b+c) 8300000 8300000 8482600
% of completion (e=b/d*100) 30% 78% 100%
Revenue to be recognised (f) 3000000 7800000 10000000
Billing done (g) 2030000 4444000 3526000
Accumulated Billing (h) 2030000 6474000 10000000
Unbilled Revenue (i = f - h) 970000 1326000 0
Cash Received (j) 1815000 3900000 4285000
Accumulated cash received (k) 1815000 5715000 10000000
Trade Receivable (l = h-k) 215000 759000 0

2) Completing the partial balance sheet:

Balance Sheet (Partial) 2018 2019
Current Assets
Accounts Receivable $215000 $759000
Construction In Progress (Unbilled Revenue) $970000 $1326000
Less: Billings ($2030000) ($4444000)
Cost and Profit in excess of Billing ($2030000) ($4444000)